OpenStudy (anonymous):

\[\sqrt{x}+y=11 and \sqrt{y}+x=7 (solve buy elimination method)\]

OpenStudy (anonymous):

Something you can try: \[\sqrt x+y=11~~\implies~~\sqrt x=11-y~~\implies~~x=(11-y)^2\] Subtracting this equation from the second gives \[(\sqrt y+x)-x=7-(11-y)^2\] This eliminates the \(x\), giving \[\sqrt y=7-(11-y)^2\]

OpenStudy (anonymous):

solve it must not by putting method

OpenStudy (anonymous):

i.e. putting value like 1,2,3

OpenStudy (anonymous):

I don't quite understand what you're trying to say. I'm using elimination, as requested by the original question.